OpenStudy (anonymous):

Is this right? Solve 12x + 2(5 – 2x) = 2x – 6(7 – x). Please show all of your work. 12x+2(5-2x) = 2x-6(7-x) 12x+10-4x = 2x-36-6x 12x+4x-(-2x)+6x=36-10 16x= 26 16x/16= 26/16 x=1 10/16 x = 1 5/8 simplest form

OpenStudy (anonymous):

2nd line is off -6*7 is not 36 - *-=+

OpenStudy (anonymous):

thanks i always forget to put the sign

OpenStudy (anonymous):

you can thank me by pushing the good answer button

OpenStudy (anonymous):

Oh, and btw, just in case you didn't notice Ty... the sign was right, the problem was that -6*7=-42, not 36.

OpenStudy (anonymous):

not a problem.

OpenStudy (anonymous):

thanks. good catch gandalfwiz

OpenStudy (anonymous):

12x+2(5-2x) = 2x-6(7-x) 12x+10-4x = 2x-42-6x 12x+4x-(-2x)+6x=-42-10 16x= -36 16x/16= -36/16 x= -2 4/16 x = -2 1/4 simplest form